# Leveraging Large Language Model for Information Retrieval-based Bug Localization

### v1

@@ V-B RQ2: Ablation Study
GenLoc 43.19 62.35 67.87 0.41 0.51
Embedding-Only 32.42 56.86 65.50 0.34 0.43
GenLoc-NoEmbed 34.28 43.29 45.71 0.31 0.38
Bug ID: 232757
Summary: Error with standard meter charts: org.eclipse.birt.chart.exception.ChartException: Could not zoomOut for step=8
Description: I am using BIRT 2.2.2. I have a report that displays several “standard meter” charts. But in some cases, some dials are not displayed and I get the following error:
org.eclipse.birt.chart.exception.ChartException: Could not zoomOut for step=8
at org.eclipse.birt.chart.factory.Generator.render(Generator.java:1287)
org.eclipse.birt.report.engine.layout.html.HTMLPageLM.layout(HTMLPageLM.java:85) org.eclipse.birt.chart.computation.withaxes.AutoScale.zoomOut(AutoScale.java:515) … 151 more. Also I have put in attachment a simplified version of a report that shows problems.
Figure 5: Bug Report from Birt Project.
The GenLoc-NoEmbed achieves only 34.61% Accuracy@1, showing that the absence of embedding guidance limits the LLM’s ability to identify relevant files effectively. Fig. 6 presents a sample bug [^4] https://bugs.eclipse.org/bugs/show_bug.cgi?id=147801 from the AspectJ project, where GenLoc successfully localized the bug, while GenLoc-NoEmbed failed. The bug report referenced a DB2 class (DB2PreparedStatement) and indicated a failure during runtime weaving [^5] Runtime weaving denotes the process of dynamically inserting additional behavior, known as aspects (e.g., logging), into a program during execution, without modifying the original source code.. Although both GenLoc and GenLoc-NoEmbed attempted to find the mentioned class or methods directly, these efforts were unsuccessful. However, GenLoc overcame this problem by invoking get_candidate_filenames(), which retrieved semantically related weaving components such as BcelClassWeaver.java and WeavingAdaptor.java. Next, by analyzing their method bodies like weave() and weaveClass(), GenLoc could identify the correct buggy file (BcelClassWeaver.java)). In contrast, GenLoc-NoEmbed lacked access to these relevant files due to its inability to retrieve semantically similar candidates, demonstrating the importance of embedding-guided retrieval in resolving non-obvious fault locations. Overall, these results highlight that embedding-based retrieval and LLM-based reasoning are complementary, and removing either degrades localization performance.
Bug ID: 147801
Summary: java.lang.ClassFormatError: Repetitive method name/signature
Description:I get a ClassFormatError trying to perform runtime-weaving on a DB2 class. The class being woven is COM.ibm.db2.jdbc.app.DB2PreparedStatement. The full message exception is: java.lang.ClassFormatError: Repetitive method name/signature in class file COM/ibm/db2/jdbc/app/DB2PreparedStatement. In examining the post-weave class file, it looks as if Aspectj is defining a duplicate method within the class called getParameterMetaData().
’COM.ibm.db2.jdbc.app.DB2PreparedStatement’ (DB2PreparedStatement.java:3971) advised by afterReturning advice from ’com.ibm.tivoli.itcam.toolkit.ai.aspectj.captureJDBC.CaptureSQLStatement’ (CaptureSQLStatement.aj:946) [with runtime test]
Figure 6: Bug Report from AspectJ Project.
VI Discussion
This paper presents the first IR-based bug localization technique that uses an LLM. As highlighted in a recent survey  [55], future research should focus on leveraging state-of-the-art LLMs such as GPT and LLaMA and exploring prompt engineering and agent-based techniques to enhance IR-based bug localization. GenLoc operationalizes this vision by integrating semantic retrieval and LLM-driven reasoning within a unified bug localization pipeline.
The results demonstrate that GenLoc outperforms existing bug localization techniques, achieving the highest Accuracy@1 and consistently ranking buggy files higher across six diverse open-source Java projects. Compared to traditional IR-based techniques such as BugLocator and BRTracer, as well as recent deep learning-based methods like DreamLoc, GenLoc delivers substantial improvements by combining embedding-based candidate retrieval with iterative reasoning powered by an LLM. Unlike BugLocator, BRTracer, and DreamLoc, GenLoc operates solely on the bug report, without relying on historical bug-fix data or change history—resources that are often unavailable in new projects. This makes GenLoc more broadly applicable to any project.
@@ VI Discussion
The average cost to localize a bug using GenLoc is \$0.012, which includes \$0.003 for embedding-based retrieval (Step-1) and \$0.008 for LLM-based reasoning (Step-2). The average processing time per bug is 48.40 seconds (8.32 seconds for Step-1 and 40.08 seconds for Step-2). Notably, this processing time aligns well with developer expectations where prior studies indicate that 90% of developers expect a bug localization technique to return results within one minute. GenLoc meets this requirement while maintaining high accuracy and low cost, making it practical for real-world debugging scenarios.
@@
Another limitation is that GenLoc may not identify all files associated with a given bug within its top-10 ranked list, which is the current output size of the system. However, a prior work from Facebook [56] suggests that predicting even a single relevant file is useful in practice since it helps route the issue to the appropriate developer or team and can serve as a starting point for discovering additional affected files through manual inspection or complementary techniques. Although this does not replace the need for comprehensive localization, it highlights the potential value of partial predictions in real-world debugging workflows.
VII Threats to Validity
@@
Internal Validity: One of the most significant threats to internal validity is the possibility of data leakage, such as bug-fixing commits being present in the LLM’s training data. However, the comparison between GenLoc and GenLoc-NoEmbed in RQ2 suggests that GenLoc’s performance is not solely attributable to model memorization. Another potential threat arises from the inherent randomness of LLMs, which may yield different results across different runs. To mitigate this, all experiments were repeated three times, and a replication package is released to support reproducibility and enable further validation by the research community. Additionally, the choice of LLM may impact the results, as different models can vary significantly in reasoning capabilities, code understanding and sensitivity to prompt structure. Hence, future evaluations should assess the robustness of GenLoc across multiple LLM variants.
Construct Validity: Construct validity concerns whether the evaluation metrics accurately measure the effectiveness of bug localization techniques. This study employs Accuracy@k, MRR@10, and MAP@10, which are widely used in prior bug localization research and considered standard for assessing ranking quality [2].
External Validity: GenLoc is evaluated on over 9,000 bug reports from six open-source Java projects of varying sizes and domains. This dataset is widely used in prior bug localization research [21]. However, the findings may still not generalize to proprietary industrial systems or projects written in other programming languages. Additional studies are needed to evaluate the effectiveness of GenLoc in broader and more heterogeneous settings.
VIII Related Work
@@
VIII-A Information Retrieval-based Bug Localization
@@
Existing IRBL approaches can be broadly categorized into two groups based on their underlying methodology. The first category [57, 12, 9, 58, 59, 47, 29] uses traditional IR models such as Latent Dirichlet Allocation (LDA) and VSM for bug localization. While Lukins et al. were the first to apply LDA for bug localization [57], BugScout extended this idea by adding bug history and file size [12]. BugLocator introduced a revised VSM incorporating file size and similar past bugs [9]. BLUiR further improved retrieval by using structured code elements such as class and method names [47], while AmalGAM combined the strengths of BugLocator and BLUiR with version history [60]. Sisman and Kak introduced bug and file modification history [58], and Kim et al.improved precision by filtering uninformative bug reports [59]. BRTracer included stack traces and code chunking [29], while later efforts such as Locus [61] and BLIA [62] incorporated fine-grained change information.
@@
The second category [63, 48, 15, 16, 17, 64, 5, 65, 66, 19, 67, 68, 21, 69, 70, 71] includes machine learning and deep learning based approaches. [63, 48] leveraged learning-to-rank model for bug localization. Early deep learning methods such as HyLoc [15], DNNLOC [16], and DeepLoc [17] used DNNs to find semantic similarity between bug report and source code. CNN-based models like BugPecker [64], DreamLoc [5], MD-CNN [65], and TRANP-CNN [66] enhanced structural understanding using ASTs, CFGs, and transfer learning. Recent approaches, including FLIM [19], FBL-BERT [67], HMCBL [68] and RLocator [21] leveraged pre-trained language models and hybrid architectures to improve semantic matching and scalability. Other models like CoLoc [69], ImbalancedBugLoc [70], and DependLoc [71] addressed specific challenges such as class imbalance, cross-project localization, and dependency-aware analysis. Unlike other approaches, this paper explores the use of large language models for IRBL.
VIII-B LLMs in Software Debugging and Automated Issue Resolution
@@
Recent studies have investigated the usage of LLMs in software debugging and automated issue resolution. A notable line of work focuses on spectrum-based bug localization, where approaches such as AutoFL [72], LLM4FL [73], and CosFL [74] leverage failing test cases to identify faulty code segments. Building on this, FuseFL [75] enhances interpretability by generating step-by-step reasoning to justify faulty lines, while FlexFL [76] further improves accuracy by combining bug reports with execution traces.
@@
In parallel, other research efforts aim to enhance bug reports using LLMs. For instance, ChatBR [77] generates missing information such as observed and expected behavior, and Plein and Bissyandé [78] investigate how LLMs can generate test cases directly from bug reports. However, none of these studies analyzed the impact of LLM in the context of IRBL, when no other information (e.g. test cases) is available.
@@
Another emerging direction involves the use of LLMs for automated issue resolution on platforms like GitHub. Techniques such as MAGIS [79], AutoCodeRover [45], SpecRover [80] and Agentless [81] aim to address software issues including bugs. However, these methods primarily focus on the repair phase, often overlooking the bug localization step, which remains a crucial and challenging part of the debugging process [82].
IX Conclusion and Future Work

### v2

@@ 5.2. RQ2: Comparison with LLM-based IRBL Techniques
GenLoc 63.36 76.85 79.65 0.59 0.69
BRaIn 22.90 51.15 61.07 0.28 0.34
LLM-BL 50.38 63.11 67.18 0.47 0.56
Agentless 28.50 52.93 65.40 0.33 0.39
Figure 5. Overlap Analysis between GenLoc and Recent LLM-based Approaches.
In Fig. 6, representative lines from a bug report [^1] https://github.com/apache/rocketmq/issues/7026 are shown where GenLoc could localize the correct bug file within top 2 rank across all trials, whereas no other technique could localize it even within top 5. The report noted that the proxy returned a null address when the master broker was down instead of failing over to a slave. GenLoc first applied embedding-based retrieval to narrow the search space, where ClusterTopicRouteService.java aligned with the bug context. It then analyzed the method signatures of this file to confirm its relevance and finally inspected method bodies to uncover the missing failover logic. In contrast, neither BRaIn nor LLM-BL could shortlist the correct file, and Agentless, despite exposing the entire repository structure to the LLM, failed because the correct file was lost in the overwhelming directory and file information. By contrast, GenLoc’s two-stage process placed the correct file within the Top-50 through embeddings and then promoted it to the Top-2 through LLM based reasoning.
Bug ID: 7027
Summary: [Bug] The proxy in the cluster mode returns null adress when master is down
Description:Describe the Bug When I connect to the clustered proxy to consume messages using the remoting protocol, the master broker happened to hang up. At this time, the address returned by the proxy was empty, which caused the request to be sent to the NameServer. Steps to Reproduce 1. Deploy tow group of brokers, such as: 1. master: broker-a, slave: broker-a-s 2. master: broker-b, slave: broker-b-s 2. Deploy a proxy 3. Create some topics in broker-a for testing, such as message produing and consuming, that is not important. 4. **Kill master broker-a**. # What Did You Expect to See? Proxy should return slave address, for consumer can continue consume message from slave. What Did You See Instead? Proxy send request to NameServe. # Additional Context _No response_
Figure 6. Bug Report from Apache RocketMQ Project.
5.3. RQ3: Ablation Study
To better understand the contribution of different components in GenLoc’s bug localization pipeline, we conduct an ablation study using the following four configurations:
  1. (1)
    GenLoc: This is the complete version of the proposed approach, where the LLM is provided with five functions, including get_candidate_filenames() to utilize embedding-based retrieval to identify potentially buggy files.
  2. (2)
    Embedding-Only: A baseline that evaluates the standalone performance of the embedding-based retrieval system, where source files are ranked solely based on their semantic similarity to the bug report, without any LLM reasoning or iterative refinement.
  3. (3)
    GenLoc-NoEmbed: A variant where the LLM operates without access to the embedding-based recommended files, i.e., the get_candidate_filenames() function is removed. The model must infer filenames directly from the bug report and continue reasoning with the remaining four functions.
  4. (4)
    GenLoc-Naive: A minimal setup, where the LLM is limited to using only two functions—search_file() and search_method()—and must rely solely on basic keyword-based exploration without access to embedding guidance or deeper code-level analysis.
Table 8. Result of Ablation Study
Technique Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
GenLoc 63.36 76.85 79.65 0.59 0.69
Embedding-Only 20.36 58.27 72.52 0.30 0.37
GenLoc-NoEmbed 46.05 56.24 57.76 0.42 0.50
GenLoc-Naive 46.31 57.25 58.78 0.43 0.50
Table 8 presents the average results on the GHRB dataset across all four configurations. We observed similar results on the subset of Ye et al. dataset, which can be found in the replication package. Results show that GenLoc achieves the best performance on all metrics, demonstrating the strength of combining embedding-based retrieval with iterative LLM analysis. The Embedding-Only variant obtains only 20.36% at Accuracy@1, indicating that retrieval alone is insufficient without deeper analysis. Similarly, the GenLoc-NoEmbed achieves only 46.05% Accuracy@1, showing that the absence of embedding guidance limits the LLM’s ability to identify relevant files. These results highlight that embedding-based retrieval and LLM-based reasoning are complementary, and removing either degrades localization performance.
The performance improvement of GenLoc over GenLoc-NoEmbed and GenLoc-Naive indicates that GenLoc goes beyond superficial cues (e.g., filenames or keywords) and performs deeper reasoning. A case [^2] https://github.com/OpenAPITools/openapi-generator/issues/19039 is shown in Fig. 7, where GenLoc-NoEmbed and GenLoc-Naive failed to localize the correct file, while GenLoc succeeded. The bug report described an error in the TypeScript Fetch generator of OpenAPI, where the modelNamePrefix option was being incorrectly applied twice in generated import statements. Misled by surface-level clues, GenLoc-NoEmbed and GenLoc-Naive attempted to locate a nonexistent file named ModelNamePrefix and also searched for the generate() method, which was unrelated to the bug and provided no useful signal. In contrast, GenLoc invoked get_candidate_filenames() to narrow down to the TypeScript-specific implementation and by analyzing method signatures and bodies in TypeScriptFetchClientCodegen.java (e.g., toModelFilename(), parseImports()), it successfully localized the buggy file.
Bug ID: 20109
Summary: [BUG] [typescript-fetch] ModelNamePrefix is added twice to import statements thereby breaking the build
Description: When generating the model files the prefix is properly added once in front of the generated files. However, in the import statements the prefix is added _twice_ resulting in erroneous statements breaking the build. When e.g. ‘SomePrefix‘ is set for ‘modelNamePrefix‘, then the following model files are generated for ‘example-for-file-naming-option.yaml‘: * ‘SomePrefixPetCategory.ts‘ * ‘SomePrefixPet.ts‘ However, _within_ ‘SomePrefixPet.ts‘ it looks like this: “‘typescript import type SomePrefixPetCategory from ’./SomePrefixSomePrefixPetCategory’; import SomePrefixPetCategoryFromJSON, SomePrefixPetCategoryFromJSONTyped, SomePrefixPetCategoryToJSON, from ’./SomePrefixSomePrefixPetCategory’; “‘ I.e. ‘SomePrefix‘ is added twice.
Figure 7. Bug Report from OpenAPI Generator Project.
Figure 8. Overlap analysis between GenLoc and Its Ablated Variants.
Although GenLoc-NoEmbed and GenLoc-Naive obtain similar results, the overlap analysis (Fig. 8) shows that they captures distinct subsets of bugs. This demonstrates the importance of combining all four functions—search_file(), search_method(), get_method_signatures_of_a_file() and get_method_body()—to achieve broader coverage. Their integration enables GenLoc to leverage both keyword-level and structure-aware reasoning.
Finally, Table 9 presents the consistency of results, measured as consensus across trials for each variant (i.e., GenLoc, GenLoc-NoEmbed and GenLoc-Naive). Here, consensus indicates the proportion of bugs that are localized in at least two or all three trials, reflecting the stability of each approach. GenLoc shows the highest stability, with 93.69% of bugs localized in at least two runs and 88.29% in all three runs. In contrast, GenLoc-NoEmbed and GenLoc-Naive attain lower values, reflecting greater variability across trials. It shows that the integration of all components not only improves accuracy but also enhances the consistency of GenLoc’s results.
Table 9. Consistency Across Trials
Consensus GenLoc (%) GenLoc-NoEmbed (%) GenLoc-Naive (%)
$\geq$ 2 trials 93.69 85.23 86.36
$\geq$ 3 trials 88.29 72.73 76.14
5.4. RQ4: Performance on Unseen Bugs
Table 10 presents the performance of GenLoc on both seen and unseen bug reports from the GHRB dataset. The results indicate that GenLoc maintains consistent performance across the two subsets. For instance, the accuracy@1 remains stable at approximately 63% for both seen and unseen bugs. To further validate this observation, we performed a Mann–Whitney U test, a non-parametric method suitable for comparing independent groups (McKnight and Najab, 2010). The results reveal no statistically significant difference in Reciprocal Rank (RR) (p-value > 0.7) and Average Precision (AP) (p-value > 0.3) between the two groups. These findings suggest that GenLoc does not rely on memorization or training data leakage and can effectively localize bugs even in reports that fall outside the temporal scope of the LLM’s training data.
Table 10. Performance on Seen and Unseen Bugs
Category Accuracy@k (%) MAP@10 MRR@10
k=1 k=5 k=10
Seen 63.41 78.05 80.90 0.60 0.70
Unseen 63.27 74.83 77.55 0.56 0.68
6. Discussion
The evaluation demonstrates that GenLoc consistently outperforms traditional IR-based, deep learning–based and recent LLM-based bug localization techniques. Its strength lies in integrating semantic retrieval with LLM-driven iterative reasoning, which ensures that relevant files are surfaced and analyzed without overwhelming the model. This design highlights the importance of context quality in LLM-driven bug localization—too little context leads to irrelevant analysis, while too much hinders reasoning. By striking this balance through embedding-based retrieval and targeted exploration, GenLoc achieves reliable and consistent performance improvements across diverse projects.
Apart from higher accuracy, GenLoc is also cost-effective. On the Ye et al. dataset, the average cost to localize a bug is \$0.010, which includes \$0.003 for embedding-based retrieval (Step 1) and \$0.007 for LLM-based reasoning (Step 2). On the more recent GHRB dataset, the cost is \$0.025 per bug, with \$0.019 for embeddings and \$0.006 for LLM reasoning. The lower per-bug cost on the Ye et al. projects, despite their larger size in terms of files and lines of code, arises from their multi-version structure: once the initial embedding is completed, only changed or newly added files need to be reprocessed in later versions. By contrast, GHRB includes more projects but far fewer bug reports per project, so a larger fraction of projects must go through the full initial embedding stage without enough subsequent bugs to offset this overhead. Nevertheless, the per-bug cost of \$0.025 remains reasonable, demonstrating that GenLoc is affordable and practical for both long-lived projects and newer projects with fewer versions. Moreover, unlike BugLocator, BRTracer and DreamLoc, GenLoc relies solely on the bug report, thereby making it broadly applicable even in projects without historical bug-fix data.
@@ 6. Discussion
In terms of efficiency, GenLoc can produce results within practical time limits. The average processing time per bug is 48.35 seconds on the Ye et al. dataset and 37.76 seconds on GHRB. These times are consistent with findings from a prior study (Kochhar et al., 2016) which reported that 90% of developers expect a bug localization technique to return results within one minute. This alignment suggests that GenLoc can operate within developer expected timeframe while also maintaining high accuracy and low cost, highlighting its potential suitability for real-world debugging scenarios.
@@
Another limitation is that GenLoc may not identify all files associated with a given bug within its top-10 ranked list, which is the current output size of the system. However, a prior work from Facebook (Pradel et al., 2020) suggests that predicting even a single relevant file is useful in practice since it helps route the issue to the appropriate developer or team and can serve as a starting point for discovering additional affected files through manual inspection or complementary techniques. Although this does not replace the need for comprehensive localization, it highlights the potential value of partial predictions in real-world debugging workflows.
7. Threats to Validity
@@
Internal Validity: One of the most significant threats to internal validity is the possibility of data leakage, such as bug-fixing commits being present in the LLM’s training data. However, results of RQ4 show that GenLoc maintains consistent performance on both seen and unseen bug reports in the GHRB dataset, indicating that GenLoc effectively localizes bugs even for reports outside the temporal scope of the LLM’s training data. Another potential threat arises from the inherent randomness of LLMs, which may yield different results across different runs. To mitigate this, all experiments were repeated three times, and a replication package is released to support reproducibility and enable further validation by the research community. Additionally, the choice of LLM may impact the results, as different models can vary significantly in reasoning capabilities, code understanding and sensitivity to prompt structure. Hence, future evaluations should assess the robustness of GenLoc across multiple LLM variants.
Construct Validity: Construct validity concerns whether the evaluation metrics accurately measure the effectiveness of bug localization techniques (Widyasari et al., 2022). This study employs Accuracy@k, MRR@10 and MAP@10, which are widely used in prior bug localization research and considered standard for assessing ranking quality (Wong and Tse, 2023; Rahman and Roy, 2018).
External Validity: GenLoc is evaluated on two complementary datasets: a large benchmark of over 9,000 bug reports from six open-source Java projects that has been widely used in prior bug localization research (Chakraborty et al., 2024), and the more recent GHRB dataset of 131 bugs from 16 projects, designed specifically for evaluating LLM-based approaches (Lee et al., 2024). While these datasets provide strong evidence of GenLoc’s effectiveness, the findings may still not generalize to proprietary industrial systems or projects written in other programming languages. Additional studies are needed to evaluate the effectiveness of GenLoc in broader and more heterogeneous settings.
8. Related Work
@@
8.1. Information Retrieval-based Bug Localization
@@
Existing IRBL approaches can be broadly categorized into three groups based on their underlying methodology. The first category (Lukins et al., 2008; Nguyen et al., 2011; Zhou et al., 2012; Sisman and Kak, 2012; Kim et al., 2013; Saha et al., 2013; Wong et al., 2014) uses traditional IR models such as Latent Dirichlet Allocation (LDA) and VSM for bug localization. While Lukins et al. were the first to apply LDA for bug localization (Lukins et al., 2008), BugScout extended this idea by adding bug history and file size (Nguyen et al., 2011). BugLocator introduced a revised VSM incorporating file size and similar past bugs (Zhou et al., 2012). BLUiR further improved retrieval by using structured code elements such as class and method names (Saha et al., 2013), while AmalGAM combined the strengths of BugLocator and BLUiR with version history (Wang and Lo, 2014). Sisman and Kak introduced bug and file modification history (Sisman and Kak, 2012), and Kim et al. improved precision by filtering uninformative bug reports (Kim et al., 2013). BRTracer included stack traces and code chunking (Wong et al., 2014), while later efforts such as Locus (Wen et al., 2016) and BLIA (Youm et al., 2017) incorporated fine-grained change information.
@@
The second category (Ye et al., 2014; Fejzer et al., 2021; Lam et al., 2015, 2017; Xiao et al., 2019; Cao et al., 2020; Qi et al., 2021; Wang et al., 2020; Iqbal et al., 2020; Liang et al., 2022; Ciborowska and Damevski, 2022; Du and Yu, 2023; Chakraborty et al., 2024; Luo et al., 2022; Anh and Luyen, 2021; Zakari et al., 2020) consists of learning-based models that apply machine learning and deep learning to capture richer semantic relationships between bug reports and source code. (Ye et al., 2014; Fejzer et al., 2021) leveraged learning-to-rank model for bug localization. Early deep learning methods such as HyLoc (Lam et al., 2015), DNNLOC (Lam et al., 2017), and DeepLoc (Xiao et al., 2019) used DNNs to find semantic similarity between bug report and source code. CNN-based models like BugPecker (Cao et al., 2020), DreamLoc (Qi et al., 2021), MD-CNN (Wang et al., 2020), and TRANP-CNN (Iqbal et al., 2020) enhanced structural understanding using ASTs, CFGs, and transfer learning. Later approaches, including FLIM (Liang et al., 2022), FBL-BERT (Ciborowska and Damevski, 2022), HMCBL (Du and Yu, 2023) and RLocator (Chakraborty et al., 2024) leveraged pre-trained language models as feature extractors to improve semantic matching and scalability. Other models like CoLoc (Luo et al., 2022), ImbalancedBugLoc (Anh and Luyen, 2021), and DependLoc (Zakari et al., 2020) addressed specific challenges such as class imbalance, cross-project localization, and dependency-aware analysis.
The third category consists of reasoning-enhanced approaches (Li et al., 2025b; Samir and Rahman, 2025; Li et al., 2025a) where language models are used as an active reasoning component. KEPT enhances the UniXcoder language model by incorporating project-specific knowledge graphs, and fine-tuning it using contrastive learning (Li et al., 2025b). BRaIn combines lexical retrieval with LLM-based method analysis and iterative refinement (Samir and Rahman, 2025), while LLM-BL employs LLM-driven query expansion and hybrid retrieval to enhance the ranking of buggy files and lines (Li et al., 2025a). GenLoc differs from prior LLM-based approaches by combining semantic retrieval with iterative code-exploration functions, allowing the LLM to selectively inspect and reason over the code base rather than relying solely on retrieval or reranking.
8.2. LLMs in Software Debugging and Automated Issue Resolution
@@
Recent studies have investigated the usage of LLMs in software debugging and automated issue resolution. A notable line of work focuses on spectrum-based bug localization, where approaches such as AutoFL (Kang et al., 2024), LLM4FL (Rafi et al., 2024), CosFL (Qin et al., 2025a) and SoapFL (Qin et al., 2025b) leverage failing test cases to identify faulty code segments. Building on this, FuseFL (Widyasari et al., 2024) enhances interpretability by generating step-by-step reasoning to justify faulty lines, while FlexFL (Xu et al., 2025) further improves accuracy by combining bug reports with execution traces.
@@
In parallel, other research efforts aim to enhance bug reports using LLMs. For instance, ChatBR (Bo et al., 2024) generates missing information such as observed and expected behavior, while (Kang et al., 2023) and (Plein and Bissyandé, 2023) investigate how LLMs can generate test cases directly from bug reports. However, none of these studies analyzed the impact of LLM in the context of IRBL, when no other information (e.g. test cases) is available.
@@
Another emerging direction involves the use of LLMs for automated issue resolution on platforms like GitHub. Techniques such as MAGIS (Tao et al., 2024), AutoCodeRover (Zhang et al., 2024), SpecRover (Ruan et al., 2025), LingmaAgent (Ma et al., 2025) and Agentless (Xia et al., 2025) aim to address software issues including bugs. However, these methods primarily focus on the repair phase, often overlooking the bug localization step, which remains a crucial and challenging part of the debugging process (Chang et al., 2025). In contrast, this paper focuses on the bug localization phase, investigating how LLMs can be leveraged to enhance IRBL.
9. Conclusion and Future Work